# Tell, Don’t Show!: Language Guidance Eases Transfer Across Domains in Images and Videos

### v1

@@ Abstract
We introduce LaGTran, a novel framework that utilizes readily available or easily acquired text descriptions to guide robust transfer of discriminative knowledge from labeled source to unlabeled target data with domain gaps. While unsupervised adaptation methods have been established to address this problem, they show limitations in handling challenging domain shifts due to their exclusive operation within the pixel-space. Motivated by our observation that semantically richer text modality has more favorable transfer properties, we devise a transfer mechanism to use a source-trained text-classifier to generate predictions on the target text descriptions, and utilize these predictions as supervision for the corresponding images. Our approach driven by language guidance is surprisingly easy and simple, yet significantly outperforms all prior approaches on challenging datasets like GeoNet and DomainNet, validating its extreme effectiveness. To further extend the scope of our study beyond images, we introduce a new benchmark to study ego-exo transfer in videos and find that our language-aided approach LaGTran yields significant gains in this highly challenging and non-trivial transfer setting. Code, models, and proposed datasets are publicly available at https://tarun005.github.io/lagtran/.
@@ 1 Introduction
We further illustrate this property in Fig. 1, where we examine the transferability of image and text classifiers trained using image or text supervision respectively between USA and Asia domains from the GeoNet dataset (Kalluri et al., 2023). We observe significantly less drop ($9.5\%$) when applying a text classifier trained on the source text to target text, compared to $17.1\%$ drop incurred when transferring an image classifier to classify target images  [^1] These observations hold for other domains with different text sources as well, as shown in the supplementary.. As text operates in a significantly lower-dimensional space, language modality naturally tends to have lesser domain gaps as opposed to images or videos. Furthermore, text descriptions often contain valuable attributes and identifiers that enhance the ability to accurately recognize images in a standard classification setting, suggesting more favorable domain robustness and discriminative properties of language descriptions compared to images.
@@
  • Demonstration of the competence of LaGTran across a variety of domain shifts, with non-trivial gains over UDA methods on challenging datasets like GeoNet (+10%), DomainNet (+3%) and the proposed Ego2Exo (+10%) datasets (Sec. 4).
@@ 2 Related Work
Our work studies the setting where it is necessary to improve accuracy on an unlabeled target domain by leveraging labels from a different source domain. A suite of methods have been proposed to study this problem using unsupervised adaptation (Ben-David et al., 2006, 2010; Ganin & Lempitsky, 2015; Long et al., 2018), where prior works propose various domain alignment strategies including MMD-based alignment (Tan et al., 2020; Long et al., 2017; Sun & Saenko, 2016; Kang et al., 2019), adversarial alignment (Bousmalis et al., 2016; Tzeng et al., 2017; Saito et al., 2017; Chen et al., 2019a; Tzeng et al., 2015; Wei et al., 2021), class-specific adaptation (Pei et al., 2018; Saito et al., 2018; Luo et al., 2019; Xie et al., 2018; Kumar et al., 2018; Gu et al., 2020), clustering (Deng et al., 2019; Park et al., 2020; Li et al., 2021a; Kalluri & Chandraker, 2022) instance-specific adaptation (Sharma et al., 2021; Kalluri et al., 2022; Wang et al., 2022), self-training (French et al., 2017; Liu et al., 2021a; Sun et al., 2022; Prabhu et al., 2022) and more recently transformers (Xu et al., 2021) or patch-based mechanisms (Zhu et al., 2023). Similar ideas have also been explored in video domain adaptation (Choi et al., 2020), with extensions to incorporate temporal alignment  (Chen et al., 2019a; Wei et al., 2023; Sahoo et al., 2021; Dasgupta et al., 2022). However, all these uda methods predominantly operate in the pixel-space, and often fall short in bridging more complicated forms of domain shift in challenging transfer settings (Kalluri et al., 2023; Prabhu et al., 2022). Alternatively, our work introduces a new paradigm to study robustness, where we build a simple framework using rich textual descriptions to overcome large domain gaps in image and video recognition tasks.
@@
Recent emergence of large-scale pre-trained vision-language foundational models such as CLIP (Radford et al., 2021) enabled strong zero-shot generalization across diverse domains and tasks (Devillers et al., 2021). However, the zeroshot inference using frozen pre-trained models still fall short of supervised fine-tuning (Radford et al., 2021; Pham et al., 2023; Andreassen et al., 2021), which in-turn suffers from poor generalization to distributions outside the fine-tuning data (Kumar et al., 2022; Wortsman et al., 2022). Prior works explored robust fine-tuning of zero-shot models, but do not leverage target domain data (Udandarao et al., 2023) or language supervision (Wortsman et al., 2022; Goyal et al., 2023) during fine-tuning. While recent works incorporate language guidance for domain generalization (Dunlap et al., 2023; Wang et al., 2024; Liu & Wang, 2023; Huang et al., 2023; Min et al., 2022), they mostly rely on domain or class descriptors and do not leverage semantically richer free form text supervision from target images during transfer. In contrast to these efforts, we show that incorporating language aided transfer yields a remarkably effective framework for improving domain robustness.
@@ 3.2 LaGTran for Cross-Domain Transfer
where $\mathcal{C}$ is the set of categories in the classification task. Using these predictions, we construct a pseudo-labeled target dataset, given by $\widehat{\mathcal{D}_{t}}=\{x_{i}^{t},\hat{y}_{i}^{t}\}_{i=1}^{N_{t}}$. Finally, we combine this pseudo-labeled target images along with manually labeled source domain images to train an image classifier backbone $\mathcal{G}$. {IEEEeqnarray}C arg min_θE_(x_i,y_i) ∼D_s L_CE (G(x_i;θ), y_i) +
@@ 4 Experiments
We first study the effectiveness of LaGTran on standard image datasets (Sec. 4.1) and extensions to open world transfer (Sec. 4.4). We then show our results on a new dataset for transfer between ego-exo views in videos (Sec. 4.5) followed by extensive ablations and insights into our framework (Sec. 4.6).
@@ 4.1 LaGTran for Image Classification
4.2 LaGTran Outperforms Prior Works on GeoNet
@@
LaGTran 63.82 63.46 58.33 56.87 60.62
Table 1: LaGTran outperforms all prior methods by ${>}10\%$ on average with the challenging GeoImnet benchmark with 600 classes and GeoPlaces with 205 classes designed for geographical transfer (Sec. 4.2). All methods use a ViT-B backbone. ${}^{\dagger}$denotes domain aware-prompting. Best values in bold, second best underlined. U:USA, A:Asia.
Source Realnormal-$\rightarrow$ Clipartnormal-$\rightarrow$ Sketchnormal-$\rightarrow$ Paintingnormal-$\rightarrow$
@@
Table 2: LaGTran sets new state-of-the-art on DomainNet-345 dataset, outperforming prior methods and baselines in most tasks. All models use Swin-B backbone, and UDA numbers are taken from (Zhu et al., 2023). ${}^{\dagger}$denotes domain aware-prompting. Best values in bold, second best underlined.
@@
We present results for GeoPlaces and GeoImnet benchmarks in Tab. 1. As noted in (Kalluri et al., 2023), previous UDA methods often fall short of bridging geographic disparities, highlighting the challenge of geographical transfer with image data alone. Notably, LaGTran achieves $60.62\%$ average Top-1 accuracy on all transfer tasks, beating all previous UDA methods and strong baselines by significant margins, providing solid validation to our transfer approach using language guidance. Specifically, LaGTran outperforms the source-only baseline by ${\sim}14\%$ and best adaptation approach PMTrans (Zhu et al., 2023) by ${\sim}10\%$ on the average accuracy, highlighting the natural benefit conferred by training while leveraging text supervision in source and target domains. LaGTran even surpasses zeroshot accuracy using domain-aware prompting on CLIP (Radford et al., 2021) by ${\sim}10\%$, while being trained on order of magnitude fewer data compared to CLIP’s hundreds of millions of image-text pairs. Remarkably, we also outperform the strongest baseline TextMatch by ${\sim}8\%$, underlining the fact that in cases when the text descriptions might not always have embedded label information directly, using labels from a source domain still has significant advantage.
4.3 LaGTran achieves new SOTA on DomainNet
We summarize the results on DomainNet in Tab. 2, where LaGTran yields large gains over several prior UDA methods and all the competitive baselines, setting new state-of-the-art on this challenging dataset. Notably, many prior methods return lesser numbers than directly training on a source model (Saito et al., 2018; Zhang et al., 2019; Du et al., 2021; Li et al., 2021a), indicating their poor scalability to natural domain shifts in large-scale data. While more recent innovations in UDA such as self-training (Sun et al., 2022) and patch-based mixing (Zhu et al., 2023), as well as zeroshot inference using CLIP offer improved performance, LaGTran still outperforms these methods on most tasks. Finally, our superior accuracy compared to both baselines TextMatch and nGramMatch, that employ target-only pseudo-labeling, underscores the significance of having access to supervised text data and labels from a source domain for enhanced target accuracy.
@@ 4.4 LaGTran Improves Transfer with Outliers
We show our results on open-world transfer setting using the GeoUniDA dataset (Kalluri et al., 2023), which examines unsupervised transfer across geographies in the presence of geographically unique classes in both source and target along with common classes. Specifically, GeoUniDA contains 62 shared classes between source and target, along with 138 private categories in each domain. We follow OVANet (Saito & Saenko, 2021) to adopt the H-score evaluation metric, which weighs equal importance to closed-set and open-set accuracies by measuring the harmonic mean of both. In addition to standard works that address outlier detection through universal adaptation (You et al., 2019; Saito & Saenko, 2021; Saito et al., 2020), we also train a baseline model using only the source domain data performing test-time outlier detection using MSP thresholding (Hendrycks & Gimpel, 2016). As shown in Tab. 3, LaGTran achieves a H-score of $61.16\%$, significantly surpassing the baseline source-only accuracy as well as all prior universal adaptation approaches by ${>}10\%$, indicating that language guidance naturally provides a strong signal to detect target samples while handling outliers in open-set target domain data.
@@ 4.5 LaGTran for Video Domain Adaptation
Despite rapid advances in methods (Chen et al., 2019b; Munro & Damen, 2020; Choi et al., 2020; Wei et al., 2023) and benchmarks (Munro & Damen, 2020; Plizzari et al., 2023) for video domain adaptation, little insight is available into their ability to address challenging settings such as transfer between ego (first-person) and exo (third-person) perspectives in videos. While prior efforts studying ego-exo transfer require paired videos from both views (Quattrocchi et al., 2023; Sigurdsson et al., 2018) or do not leverage unlabeled data in the target (Li et al., 2021b; Ohkawa et al., 2023; Xue & Grauman, 2023), limited works looked into the aspect of unsupervised domain transfer from ego to exo views due to the lack of a suitable benchmark.
Therefore, we introduce a new benchmark called Ego2Exo to study transfer between the ego and exo views in videos. We curate our dataset using the recently proposed Ego-Exo4D (Grauman et al., 2023), utilizing their keystep annotations for action labels, and atomic descriptions as text supervision. We manually remap the labels to a coarser hierarchy to ease the difficult task of predicting very fine-grained action classes from short segments (eg: add coffee beans vs. add coffee grounds). The complete details about our dataset creation process are included in the supplementary, Supp. B.
Our proposed Ego2Exo consists of 10,102 video segments labeled with actions from one of the 22 keysteps, and we split these video segments into two equal groups classwise, and collect ego-videos from one group and exo-videos from the other to create our adaptation benchmark. We finally obtain 5247 ego-videos and 4855 exo-videos capturing mutually exclusive actions and scenes. The atomic action descriptions from all the timestamps within each segment, whenever available, form the text supervision for that segment. The same procedure applied to the validation videos yields 3147 segments with both ego and exo views. The distribution of the duration of segments in the benchmark, along with the label distribution for ego and exo domains is presented in Fig. 4. The videos, labels along with the text descriptions are publicly available on our project page.
(a) Segment Duration
@@
Source Only 17.06 29.36 23.21
CDAN (Long et al., 2018) 14.08 17.03 15.56
UAR (Choi et al., 2020) 14.65 18.75 16.70
TA3N (Chen et al., 2019b) 17.25 19.54 18.40
TransVAE (Wei et al., 2023) 11.31 16.01 13.66
@@
EgoVLP (Lin et al., 2022) 2.25 14.78 8.51
LaVILA (Zhao et al., 2023) 8.32 19.99 14.15
TextMatch 13.98 15.92 14.95
nGramMatch 14.96 21.51 18.24
LaGTran 22.75 43.53 33.14
Target Sup. 28.91 49.88 39.40
@@
Firstly, we highlight the importance of studying robustness across ego and exo views in Tab. 4 by examining the ego-test accuracy of a model trained directly on ego videos, which achieves 49.88%, compared to a model transferred from exo-videos, which only achieves 29.36%. Similarly, a model trained on ego videos achieve only 17.06% for recognition in exo view, compared to a potential 28.91% achievable by training directly on exo videos, indicating a significant domain shift. Current state-of-the-art video adaptation methods (Wei et al., 2023) yield limited gains to bridge these gaps, highlighting the need for novel approaches to address this challenge. Moreover, zeroshot video classification accuracy using EgoVLP (Lin et al., 2022) and LaVILA (Zhao et al., 2023) also show limited gains. Notably, LaGTran which efficiently leverages action descriptions available alongside the videos, achieves an accuracy of 33.14% on average significantly outperforming the source-only baseline by $10\%$ and prior adaptation methods by ${>}15\%$. LaGTran also outperforms pseudo-labeling using nGramMatch or TextMatch, as the text descriptions, independently developed from keystep labels, often lack utility for deciphering the action category labels on their own. We however note the substantial scope for further improvement in future, both in terms of the low within-domain accuracy as well as the remaining gap to supervised target accuracy.

### v3

@@ Abstract
We introduce LaGTran, a novel framework that utilizes text supervision to guide robust transfer of discriminative knowledge from labeled source to unlabeled target data with domain gaps. While unsupervised adaptation methods have been established to address this problem, they show limitations in handling challenging domain shifts due to their exclusive operation within the pixel-space. Motivated by our observation that semantically richer text modality has more favorable transfer properties, we devise a transfer mechanism to use a source-trained text-classifier to generate predictions on the target text descriptions, and utilize these predictions as supervision for the corresponding images. Our approach driven by language guidance is surprisingly easy and simple, yet significantly outperforms all prior approaches on challenging datasets like GeoNet and DomainNet, validating its extreme effectiveness. To further extend the scope of our study beyond images, we introduce a new benchmark called Ego2Exo to study ego-exo transfer in videos and find that our language-aided approach LaGTran yields significant gains in this highly challenging and non-trivial transfer setting. Code, models and proposed datasets are publicly available at https://tarun005.github.io/lagtran/.
@@ 1 Introduction
We further illustrate this property in Fig. 1, where we examine the transferability of image and text classifiers trained using image or text supervision respectively between USA and Asia domains from the GeoNet dataset (Kalluri et al., 2023). We observe significantly less drop ($9.5\%$) when applying a text classifier trained on the source text to target text, compared to $17.1\%$ drop incurred when transferring an image classifier to target images. As text operates in a significantly lower-dimensional space, language modality naturally has lesser domain gaps as opposed to images or videos. Furthermore, text descriptions often contain valuable attributes and identifiers that enhance the ability to accurately recognize images in a standard classification setting, suggesting more favorable domain robustness and discriminative properties of language descriptions compared to images.
@@
  • Demonstration of the competence of LaGTran across a variety of domain shifts, with non-trivial gains over UDA methods on challenging datasets like GeoNet (+10%), DomainNet (+3%) and the proposed Ego2Exo (+4%) datasets (Sec. 4).
@@ 2 Related Work
A suite of methods have been proposed to improve accuracy on an unlabeled target domain by leveraging labels from a different source domain using unsupervised adaptation (Ben-David et al., 2006, 2010; Ganin & Lempitsky, 2015; Long et al., 2018), where prior works propose various domain alignment strategies including MMD-based (Tan et al., 2020; Long et al., 2017; Sun & Saenko, 2016; Kang et al., 2019), adversarial (Bousmalis et al., 2016; Tzeng et al., 2017; Saito et al., 2017; Chen et al., 2019a; Tzeng et al., 2015; Wei et al., 2021), class-specific adaptation (Pei et al., 2018; Saito et al., 2018; Luo et al., 2019; Xie et al., 2018; Kumar et al., 2018; Gu et al., 2020), clustering (Deng et al., 2019; Park et al., 2020; Li et al., 2021a; Kalluri & Chandraker, 2022) instance-specific adaptation (Sharma et al., 2021; Kalluri et al., 2022; Wang et al., 2022), self-training (French et al., 2017; Liu et al., 2021a; Sun et al., 2022; Prabhu et al., 2022) and more recently transformers (Xu et al., 2021) or patch-based mechanisms (Zhu et al., 2023). Similar ideas have also been explored in video domain adaptation (Choi et al., 2020), with extensions to incorporate temporal alignment  (Chen et al., 2019a; Wei et al., 2023; Sahoo et al., 2021; Dasgupta et al., 2022). However, all these uda methods predominantly operate in the pixel-space, and often fall short in bridging more complicated forms of domain shift in challenging transfer settings (Kalluri et al., 2023; Prabhu et al., 2022). While some contemporary efforts utilize pre-trained CLIP models (Lai et al., 2023, 2024; Zhang et al., 2023) or LLMs (Chen et al., 2024) for domain alignment, we show that per sample natural language guidance can be equally effective. Based on this, our work introduces a new paradigm to study robustness, where we build a simple framework using rich textual descriptions to overcome large domain gaps in image and video recognition tasks.
@@
Recent emergence of large-scale pre-trained vision-language foundational models such as CLIP (Radford et al., 2021) enabled strong zero-shot generalization across diverse domains and tasks (Devillers et al., 2021). However, the zeroshot inference using frozen pre-trained models still fall short of supervised fine-tuning (Radford et al., 2021; Pham et al., 2023; Andreassen et al., 2021), which in-turn suffers from poor generalization to distributions outside the fine-tuning data (Kumar et al., 2022; Wortsman et al., 2022). Prior works explored robust fine-tuning of zero-shot models, but do not leverage target domain data (Udandarao et al., 2023) or language supervision (Wortsman et al., 2022) during fine-tuning. While recent works incorporate language guidance for domain generalization (Dunlap et al., 2023; Wang et al., 2024; Liu & Wang, 2023; Huang et al., 2023; Min et al., 2022; Gokhale et al., 2021), they mostly rely on domain or class descriptors and do not leverage semantically richer free form text supervision from target images during transfer. One work which is closest to ours is (Goyal et al., 2023), which uses label names as text descriptions while we use free-form captions for images or atomic annotations for actions. In video recognition literature, prior works seek to align ego and exo views using language pre-training (Xu et al., 2024; Huang et al., 2024). In contrast to these efforts, we show that incorporating language aided transfer through diverse supervision yields a remarkably effective framework for improving domain robustness in both images and videos.
@@ 3.2 LaGTran for Cross-Domain Transfer
where $\mathcal{C}$ is the set of categories in the classification task. Using these predictions, we construct a pseudo-labeled target dataset, given by $\widehat{\mathcal{D}_{t}}=\{x_{i}^{t},\hat{y}_{i}^{t}\}_{i=1}^{N_{t}}$. Finally, we combine this pseudo-labeled target images along with manually labeled source domain images to train an image classifier backbone $\mathcal{G}$ by sampling an equal number of images from both source and target in each mini-batch to eliminate effects caused by different dataset sizes. {IEEEeqnarray}C arg min_θE_(x_i,y_i) ∼D_s L_CE (G(x_i;θ), y_i) +
@@ 4.1 LaGTran for Image Classification
LaGTran Outperforms UDA on GeoNet
@@
LaGTran 63.67 64.16 56.14 57.02 60.24
Table 1: LaGTran outperforms all prior methods by ${>}10\%$ on average with the challenging GeoImnet benchmark with 600 classes and GeoPlaces with 205 classes designed for geographical transfer. All methods use a ViT-B backbone. denotes domain aware-prompting. Best values in bold, second best underlined. U:USA, A:Asia.
[…]
@@
Table 2: LaGTran sets new state-of-the-art on DomainNet-345 dataset, outperforming prior methods and baselines in most tasks. All models use Swin-B backbone, and UDA numbers are taken from (Zhu et al., 2023). denotes domain aware-prompting. Best values in bold, second best underlined.
@@
As noted in (Kalluri et al., 2023), previous UDA methods often fall short of bridging geographic disparities, highlighting the challenge of geographical transfer with image data alone. From Tab. 1, LaGTran achieves $60.24\%$ average Top-1 average accuracy on GeoNet, beating all previous UDA methods and strong baselines by significant margins, providing solid validation to our transfer approach using language guidance. Specifically, LaGTran outperforms the source-only baseline by ${\sim}14\%$ and best adaptation approach PMTrans (Zhu et al., 2023) by ${\sim}10\%$ on the average accuracy, highlighting the natural benefit conferred by training while leveraging text supervision in source and target domains. LaGTran even surpasses zeroshot accuracy using domain-aware prompting on CLIP (Radford et al., 2021) by ${\sim}10\%$, while being trained on order of magnitude fewer data compared to CLIP’s hundreds of millions of image-text pairs. Remarkably, we also outperform the strongest baseline TextMatch by ${\sim}8\%$, underlining the fact that in cases when the text descriptions might not always have embedded label information directly, using labels from a source is still advantageous.
LaGTran achieves new SOTA on DomainNet
We summarize the results on DomainNet in Tab. 2, where LaGTran yields large gains over several prior UDA methods and all the competitive baselines, setting new state-of-the-art. Notably, many prior methods return lesser numbers than directly training on a source model (Saito et al., 2018; Zhang et al., 2019; Du et al., 2021; Li et al., 2021a), indicating their poor scalability to natural domain shifts in large-scale data. While more recent innovations in UDA such as self-training (Sun et al., 2022) and patch-based mixing (Zhu et al., 2023) offer improved performance, LaGTran still outperforms these methods on most tasks. Finally, our superior accuracy compared to both baselines TextMatch and nGramMatch, that employ target-only pseudo-labeling, underscores the significance of having access to supervised text data and labels from a source domain for enhanced target accuracy. Further, a potential explanation for limited benefits being observed in LaGTran, as well as all previous UDA methods, when compared to CLIP under non-real to real transfer could be the precise match between images from standard categories in the real-target domain and the multi-million-scale training data utilized in CLIP. This alignment potentially eliminates any significant domain shift between the train and test settings, unlike in LaGTran where we train on non-real images yet achieve accuracies that are competitive to CLIP. Notably, LaGTran still outperforms CLIP on all domains in GeoNet (+10%) and most domains in DomainNet (upto +7%) while being trained on multiple-orders of magnitude lesser image-text pairs than CLIP.
@@ 4.2 LaGTran Improves Transfer with Outliers
We show our results on open-world transfer setting using the GeoUniDA dataset (Kalluri et al., 2023), which examines unsupervised transfer across geographies in the presence of geographically unique classes in both source and target along with common classes. Specifically, GeoUniDA contains 62 shared classes between source and target, along with 138 private categories in each domain. We follow OVANet (Saito & Saenko, 2021) to adopt the H-score evaluation metric, which gives equal importance to closed-set and open-set accuracies by measuring the harmonic mean of both. In addition to standard works that address outlier detection through universal adaptation (You et al., 2019; Saito & Saenko, 2021; Saito et al., 2020), we also train a baseline model using only the source domain data performing test-time outlier detection using MSP thresholding (Hendrycks & Gimpel, 2016). As shown in Tab. 3, LaGTran achieves a H-score of $61.16\%$, significantly surpassing the baseline source-only accuracy as well as all prior universal adaptation approaches by ${>}10\%$, indicating that language guidance naturally provides a strong signal to detect target samples while handling outliers in open-set target domain data.
@@ 4.3 LaGTran for Video Domain Adaptation
Despite rapid advances in methods (Chen et al., 2019b; Munro & Damen, 2020; Choi et al., 2020; Wei et al., 2023) and benchmarks (Munro & Damen, 2020; Plizzari et al., 2023) for video domain adaptation, little insight is available into their ability to address challenging settings such as transfer between ego (first-person) and exo (third-person) perspectives in videos. While prior efforts studying ego-exo transfer require paired videos from both views (Quattrocchi et al., 2023; Sigurdsson et al., 2018; Huang et al., 2024) or do not leverage target unlabeled data (Li et al., 2021b; Ohkawa et al., 2023), limited works study unsupervised domain transfer from ego to exo views due to the lack of a suitable benchmark.
Therefore, we introduce a new benchmark called Ego2Exo to study transfer between the ego and exo views in videos. We curate our dataset using the recently proposed Ego-Exo4D (Grauman et al., 2023), utilizing their keystep annotations for action labels, and atomic descriptions as text supervision. We manually remap the labels to a coarser hierarchy to ease the difficult task of predicting very fine-grained action classes from short clips (eg: add coffee beans vs. add coffee grounds). Complete details about our dataset creation process are included in Supp. B. Our proposed Ego2Exo consists of video segments labeled with actions from one of the 24 keysteps from ego and exo views of the corresponding actions. We obtain 4100 ego-videos and 4986 exo-videos capturing variety of actions and scenes. The atomic action descriptions from all the timestamps within each segment form the text supervision for that segment. The same procedure is applied to the validation videos yielding 3147 segments with both ego and exo views. The distribution of the duration of segments in the benchmark, along with the label distribution for ego and exo domains is presented in Fig. 4. We provide more details about the construction of the dataset in the supplementary material. The videos, labels along with the text descriptions are publicly available on our project page.
@@
Source Only 8.39 15.66 12.03
TA3N (Chen et al., 2019b) 6.92 27.95 17.44
TransVAE (Wei et al., 2023) 12.06 23.34 17.70
@@
EgoVLP (Lin et al., 2022) 5.89 19.35 12.62
LaVILA (Zhao et al., 2023) 5.86 23.16 14.51
TextMatch 10.36 13.57 11.97
nGramMatch 11.50 15.46 13.98
LaGTran 12.34 30.76 21.55
Target Sup. 17.91 33.19 25.55
@@
Firstly, we highlight the importance of studying robustness across ego and exo views in Tab. 4 by examining the ego-test accuracy of a model trained directly on ego videos, which achieves 33.19%, compared to a model transferred from exo-videos, which only achieves 15.66%. Similarly, a model trained on ego videos achieve only 8.4% for recognition in exo view, compared to a potential 17.91% achievable by training directly on exo videos, indicating a significant domain shift. Current state-of-the-art video adaptation methods (Wei et al., 2023) yield limited gains to bridge these gaps, highlighting the need for novel approaches to address this challenge. Moreover, zeroshot video classification accuracy using EgoVLP (Lin et al., 2022) and LaVILA (Zhao et al., 2023) also show limited gains. Notably, LaGTran which efficiently leverages action descriptions available alongside the videos, achieves an accuracy of 21.55% on average significantly outperforming the source-only baseline by $9\%$ and prior adaptation methods by ${>}4\%$. LaGTran also outperforms pseudo-labeling using nGramMatch or TextMatch, as the text descriptions, independently developed from keystep labels, often lack utility for deciphering the action category labels on their own. We also note the substantial scope for further improvement in future, both in terms of the low within-domain accuracy as well as the remaining gap to supervised target accuracy.